OpenStudy (anonymous):

solve 8x 3 - 12x2 -20x = o

OpenStudy (anonymous):

24-24-20x=0 x=0 ???

OpenStudy (anonymous):

is it cube there

OpenStudy (anonymous):

4x(2x^2-3x-5)=0 x(x+1)(2x-5)=0 x=0,-1,5,2.5